OpenStudy (anonymous):

Use composite of functions to determine whether f and g are inverses of one another. f(x)=x^3+2; g(x)=cubed rt x-2 I am really having difficulties grasping this concept.

OpenStudy (anonymous):

You have to show that \(f(g(x))=x\) and \(g(f(x))=x\). Do you understand what a composite function is?

OpenStudy (anonymous):

yes, the function of a function. which is f(g(x)=cubed rt (x-2)^2+2

OpenStudy (anonymous):

Your notion of a composition is right, but \[f(g(x))=f(\sqrt[3]{x-2})=\left(\sqrt[3]{x-2}\right)^3+2\] Simplifying that, you should have it equal to \(x\).

OpenStudy (anonymous):

thanks ... I did squared the cubed rt , instead of cubing it. Thanks.

OpenStudy (anonymous):

You're welcome. The work for \(g(f(x))\) is pretty similar, I'll leave it to you.